# Tailoring Polarization in WSe Quantum Emitters through Deterministic Strain Engineering

### 16 Feb 2024 v1

@@ II.2 Novel nanopillar designs
Figure 4b presents the corresponding PL image of the structure, showing the increased photon emission between the nanowrinkle (Q1). Furthermore, increased photon emission is also observed in the left corner of the top triangle, which is attributed to a nanowrinkle formed along the edge of the structure. Evidently, the separation between the two regions is sufficient to characterize them individually.
Additionally, Figure S5 presents four different SEM images of exemplary bowtie structures, all hosting a nanowrinkle connecting the two triangles, with a bar graph illustrating the general trend of directionality. Even though a statistical analysis is not possible in such a small sample, the trend is obvious, as the ’misaligned’ nanowrinkles do not deviate by more than three degrees.
After examining all suggested structural designs, we have demonstrated significant control over the polarization of the produced emitters and their spatial isolation. These findings provide crucial insights for the future integration of these quantum emitters into a specific mode of desired photonic structures, e.g., the TE mode of a planar ridge waveguide. Understanding key aspects of deterministic polarization control of quantum emitters and fabrication limitations paves the way for their effective utilization as functional single-photon sources in quantum photonics.
@@ II.3 Characterization of single-photon emission
In order to assess the quality of the fabricated SPEs based on the novel structures presented in this work, we proceed to the optical characterization of an exemplary WSe nanowrinkle quantum emitter. To perform µPL measurements, we employ a 532 nm femtosecond pulsed laser (80 MHz) to optically excite the emitter. Figure 5a shows a broad range (600-900 nm) µPL spectrum of an exemplary WSe nanowrinkle emitter with a central wavelength of 806.1 nm. A 750 nm long-pass filter was used to spectrally isolate the single-photon emission line from the delocalized exciton emission and excitation laser. The inset of Figure 5a shows a high-resolution PL spectrum of the same emitter around its central wavelength, which exhibits a prominent zero-phonon line (ZPL) with a lower-energy phonon side band (PSB). The spectrum is fitted using a sum of a Lorentzian, representing the ZPL contribution (dashed red line), and a Gaussian, which captures the information related to the PSB contribution (dashed green line). We extract the linewidths of the two fitted curves, giving values of nm and nm, respectively, with the central wavelength of the PSB being red-detuned by 0.6 nm from the ZPL, resulting in an asymmetric PL emission spectrum. The increased contribution of the PSB to the total emission, evidenced by the ratio between the intensities ZPL : PSB = 49.2 % : 50.8 %, suggests a strong exciton-phonon coupling for WSe monolayer SPEs. Mitryakhin et al. (2023) Even though the linewidth of the emission line is not Fourier-limited, being able to attain a narrow ZPL and resolve the PSB is attributed to the fabrication method that results in improved spatial isolation and a high linearly polarized emission. In addition, the spectral diffusion of the emission lines can also be affected under different excitation schemes. The excitation of the monolayer by a CW LED introduces thermal fluctuation and severe charge noise around the emitter due to photo-excited free carriers, leading to additional broadening of the emission line. To visualize the effect of additional broadening under strong CW LED excitation, we measured PL spectra under CW 470 nm LED and 532 nm pulsed laser excitation from the same emitter (see Supporting Note 5 and Figure S6). Indeed, the obtained PL spectrum under CW LED excitation exhibits a much broader and red-shifted emission line, in contrast to the pulsed excitation. To obtain the decay time of the exemplary emitter, we perform a time-resolved PL measurement. The result is presented in Figure 5b. Two different decay dynamics are clearly visible from the plot and are fitted with two separate exponential functions. The faster one exhibits a radiative lifetime at ns while the slow one is at ns. The slow decay (purple fit in 5b) is the one that dominates the time-resolved response, making the greatest contribution to the overall curve, and can be credited to the single-photon emission. Moreover, the lifetime ranges reported from previous works with regard to localized SPEs on WSe monolayer on dielectric substrates Kumar et al. (2015); Tripathi et al. (2018); So et al. (2021); Iff et al. (2021); Branny et al. (2017); He et al. (2015); Srivastava et al. (2015) agree with the reported results. The fast decay may be attributed to the instrument response function (IRF), while the abnormal increase in counts before the second decay begins may be caused by several factors. Trap states or excitation transfer processes could lead to delayed decay and an initial rise in lifetime. Concurrently, non-radiative pathways, potentially involving interactions with defects, may compete with radiative decay, contributing to this rise. Additionally, exciton dynamics, including formation and migration processes, may also influence the lifetime dynamics.
To study the stability of the emission signal from the fabricated WSe nanowrinkle, we obtain the high-resolution time trace of the PL emission with an integration time of 0.2 seconds per frame under the same above-band laser excitation scheme. We observe some wobbling of the emission line, which could attributed to the unstable charge environment and noise due to photo-excited free carriers under the above-band excitation. For a qualitative analysis of this fluctuation, we perform a fitting on the corresponding spectra, which reveals a fluctuation of 0.5 meV and an average central position of 1.5365 eV. This limitation needs to be addressed, as it could act as a bottleneck for indistinguishability as one of the potential sources of unwanted inhomogeneous broadening. To overcome the problem, several strategies could be adopted, for instance, the use of an encapsulation agent (e.g., hBN Parto et al. (2021)), electrostatic gating Lenferink et al. (2022), or employing advanced excitation schemes, i.e., resonant and SUPER scheme Karli et al. (2022). Moreover, in our out-of-plane electric contacts could also be employed to stabilize the charge environment and simultaneously provide a degree of tunability of the source via the Stark effect Noh et al. (2018).
@@
To assess the purity of the exemplary SPE, we performed the second-order correlation measurement under the aforementioned 80 MHz pulsed excitation scheme, using a 750 nm LPF to isolate the main emission line and a typical HBT setup (Figure 5d). We observe an overlap of the peaks at non-zero delay times, separated by the laser repetition time of 12.5 ns, resulting from the increased lifetime of the emitter. More importantly, the significantly diminished coincidences at zero time delay indicate the high purity of single photons emitted by the SPE. To quantify this figure of merit, we employ a double-sided exponential function fitting for the adjacent peaks, which yields a value of . This result demonstrates the exceptional single-photon purity that can be achieved through the proposed fabrication novelties on WSe monolayer quantum emitters.
Given the showcased control over the degree of polarization of single photons and spatial localization of the emitters, future works could integrate the SPEs into nanophotonic cavities to enhance the radiative emission rate via the Purcell effect. Decreasing the spontaneous emission lifetime of our nanowrinkle emitters would contribute to both increasing the photon indistinguishability Liu et al. (2018) and efficiently coupling the emitted photons to the optical cavity mode Gregersen et al. (2013). Moreover, in-depth studies on the injection of electric and magnetic fields could contribute to a more comprehensive understanding of the fundamental mechanisms governing single-photon emission in these materials and provide us with spectrally stable sources by minimizing charge noise. This could lead to the introduction of tunability over key characteristics of the emitters (e.g., main emission wavelength). Theoretical investigations may focus on exciton-phonon interactions, providing crucial insights into the advancement of SPEs in TMD platforms. At the same time, advanced excitation schemes like the SUPER scheme Bracht et al. (2021); Karli et al. (2022) could potentially minimize the interaction with phonons, resulting in a narrower ZPL.
@@ III Conclusion
In conclusion, we presented an efficient method for generating highly polarized single photons from orientation-controlled nanowrinkles in monolayer WSe. This method employs novel nanostructures to deterministically induce one-directional strain, enabling precise control of the polarization of emitted single photons. We have successfully demonstrated single photon emission from WSe nanowrinkle emitters with high purity () and a high degree of polarization (99 4%). These findings provide crucial insights for the future integration of these quantum emitters into a specific mode of desired photonic structures, e.g., the TE mode of a planar ridge waveguide. Understanding the fabrication limitations and critical aspects of deterministic polarization control of quantum emitters are essential steps toward their effective utilization as functional single-photon sources in quantum photonics.
@@ IV.1 Sample preparation
Nanopillar fabrication: The substrate was prepared by thermally oxidizing 110 nm of SiO on a silicon wafer. Alignment marks were patterned by UV lithography (using the AZ 5214E positive resist) and subsequent electron-beam evaporation of 5 nm of Ti and 50 nm of Au. After cleaving the wafer into chips, the fabrication of the nanopillars was carried out by spin coating a high-resolution negative e-beam resist (hydrogen silsesquioxane or HSQ, XR-1541-006) at 3000 rpm for 1 min, followed by two soft baking steps at 120 C and 220 C, both for 2 minutes. The resist was patterned by an e-beam writer (JEOL 9500, 100 kV, 6 nA) with a dose of 11000 µC/cm and developed in a 1:3 solution of AZ 400K : HO to obtain 150 nm-tall glass nanopillars in four different shapes, namely cylindrical pillars, three-pointed stars, five-pointed stars, and bowties. It is worth remarking that a high dose is required to cure the spin-on dielectric, whose composition becomes SiO. Other negative e-beam resists with high sensitivity (activated at doses as low as 30 µC/cm) have also been tested (Supporting Note S5).
@@
Defect fabrication: Defects were introduced on the monolayer lattice by bombarding the strained areas of the material with an electron beam (JEOL 9500 e-beam writer, 100 kV, 1000 µC/cm).
@@ IV.2 Optical characterization
Photoluminescence measurements: The PL measurements were obtained with a custom-built low-temperature micro-photoluminescence (µPL) setup. The sample is mounted on a closed-cycle cryostat (attoDRY800) operating at a base temperature of 4 K. The cryostat is equipped with piezoelectric nanopositioners and a low-temperature microscope objective ( NA = 0.82) located inside the cryostat. The sample was excited with an LED at 470 nm for the continuous-wave (CW) excitation and an 80 MHz femtosecond laser (Chameleon Ultra II, Coherent) at 550 nm for the pulsed excitation. The PL spectra were acquired with a fiber-coupled spectrometer (iHR 550, Horiba). A polarizer mounted on a motorized rotation mount was placed on the collection path for polarization-resolved measurements. PL images were acquired with an integration time of 1 s using a CMOS camera (Tucsen Dhyana, 4 MP).
@@ IV.3 Sample imaging
Figure S1: Schematic representation of the fabrication flow, Figure S2: Comparison of PL spectra before and after e-beam irradiation, Figure S3: Investigation on the morphological profile of the nanowrinkles, Figure S4: SEM images of TS and FS structures with a statistical analysis on the misalignment of the nanowrinkles, Figure S5: SEM images of BT structures with a histogram of misalignment of the nanowrinkles, Figure S6: Comparison of the PL emission of an exemplary quantum emitter under two excitation schemes (LED, 470 nm and pulsed laser, 532 nm), Figure S7: AFM image of a WSe monolayer flake on a TS structure made from AR-N 7520.11.
@@
The authors acknowledge support from the European Research Council (ERC-StG “TuneTMD”, grant no. 101076437) and the Villum Foundation (grant no. VIL53033). The authors also acknowledge the European Research Council (ERC-CoG“Unity”, grant no.865230). NG acknowledges support from the European Union’s Horizon 2020 Research and Innovation Program under the Marie Skolodowska-Curie Grant Agreement no. 861097.

### 20 Mar 2024 v2

@@ II.2 Control of nanowrinkle orientation via novel nanopillar designs
Figure 4b presents the corresponding PL image of the structure, showing increased photon emission in the nanowrinkle (Q1). Furthermore, increased photon emission is also observed in the left corner of the top triangle, which is attributed to a nanowrinkle formed along the edge of the structure. Nevertheless, the separation between the two regions is sufficient to characterize them individually. Figure 4c presents the polarization-resolved µPL spectra, with the extracted intensities reported in the polar plot of Figure 4d. Once again, the orientation of the in-plane polarization agrees with the directionality of the nanowrinkle, with an extracted DOLP of 92 .
In addition, Figure S7 presents SEM images of four different exemplary BT structures, all hosting a nanowrinkle connecting the two triangles, with a bar graph illustrating the general trend of directionality. Most BT nanowrinkles are aligned along within from the desired orientation. In contrast to the TS and FS designs, the BT seems to address all the roadblocks, achieving better alignment of the nanowrinkle while sufficiently localizing the emitters and isolating them from the surrounding environment.
After examining all suggested designs, we have demonstrated BT structures that effectively control the in-plane polarization of the emitted single photons while dictating the spatial localization of the SPEs.
@@ II.3 Characterization of single-photon emission
In order to assess the quality of the fabricated SPEs based on the novel structures presented in this work, we proceed to the optical characterization of an exemplary WSe nanowrinkle quantum emitter. For µPL measurements, we alter the excitation scheme from the CW LED to a 532 nm femtosecond pulsed laser (80 MHz) to optically excite the emitter. Supporting Note S3 and Figure S8 of the Supporting Information present the reasoning for that change. Figure 5a shows a broad range (600-900 nm) µPL spectrum of an exemplary WSe nanowrinkle emitter with a central wavelength of 806.1 nm. A 750 nm long-pass filter was used to spectrally isolate the single-photon emission line from the delocalized exciton emission and excitation laser. The inset of Figure 5a shows a high-resolution PL spectrum of the same emitter around its central wavelength, which exhibits a prominent zero-phonon line (ZPL) with a lower-energy phonon side band (PSB). The spectrum is fitted using a sum of a Lorentzian, representing the ZPL contribution (dashed red line), and a Gaussian, which captures the information related to the PSB contribution (dashed green line). We extract the linewidths of the two fitted curves, giving values of nm and nm, respectively, with the central wavelength of the PSB being red-detuned by 0.6 nm from the ZPL, resulting in an asymmetric PL emission spectrum. The increased contribution of the PSB to the total emission, evidenced by the ratio between the intensities ZPL : PSB = 49.2 % : 50.8 %, suggests a strong exciton-phonon coupling for WSe monolayer SPEs. Mitryakhin et al. (2023) Even though the linewidth of the emission line is not Fourier-limited, being able to attain a narrow ZPL and resolve the PSB is attributed to the fabrication method that results in improved spatial isolation and a high linearly polarized emission.
To investigate the decay dynamics of the emitter, we perform the time-resolved PL (TRPL) measurement under the above-band pulsed excitation (Figure 5b). Our data reveal two distinct decays, which are fitted with two separate exponential functions to extract the corresponding decay rates. The origin of the initial fast decay ( ns) is not clear at this stage; however, it seems to come from a rapid process, limited by the instrument response function (IRF). A more detailed analysis is presented in Supplementary Note S5 and Figure S10 of the Supporting Information.
Nevertheless, despite its increased amplitude, the contribution of the fast decay accounts only for of the total curve. In contrast, the slow decay ( ns) is the one that dominates the time-resolved response, making the largest contribution to the overall curve. Moreover, the calculated is within the range of the reported lifetimes for localized SPEs in WSe monolayer on dielectric substratesKumar et al. (2015); Tripathi et al. (2018); So et al. (2021); Iff et al. (2021); Branny et al. (2017); He et al. (2015); Srivastava et al. (2015). Finally, the abnormal increase in counts before the second decay begins may be caused by several factors. We attribute this rising behavior to trap states or excitation transfer processes that could lead to delayed decay and an initial rise in lifetime. Additionally, by potentially involving interactions with defects, non-radiative pathways may compete with the radiative decay, contributing to this rise.
While these complex decay dynamics require further investigation, their spectral stability is another critical aspect influencing the overall quality of SPEs. To study the stability of the emission signal from the fabricated WSe nanowrinkle, we obtain the high-resolution time trace of the PL emission with an integration time of 0.2 seconds per frame under the same above-band laser excitation scheme. We observe wobbling of the emission line, which could be attributed to the unstable charge environment and noise due to photo-excited free carriers under the above-band excitation. This unstable environment can be explained by the lack of encapsulation with a dielectric (e.g., hBN), which could minimize the interactions between the free charges and our SPE. For a qualitative analysis of this fluctuation, we perform a Gaussian fitting on the corresponding histogram presented in Figure S11, which reveals a fluctuation of the ZPL over 0.5 meV around the mean central position of 1.5365 eV.
@@
To evaluate the single-photon purity from the fabricated emitters, we performed the second-order correlation measurement under the aforementioned 80 MHz pulsed excitation using a typical HBT setup (Figure 5d). The emission line from a representative nanowrinkle emitter was spectrally isolated by implementing a 750 nm long-pass filter. We observed peaks overlap at non-zero delays, separated by the laser repetition time of 12.5 ns, resulting from the increased lifetime of the emitter. More importantly, the significantly low coincidence counts at zero time delay indicate the high purity of single-photon emission from the sample. To quantitatively assess single-photon purity, we employed a double-sided exponential function fitting for the adjacent peaks, yielding a value of .
The results presented in the previous sections demonstrate the effectiveness of the method proposed in this study. Our novel approach reliably produces high-quality single-photon emitters characterized by both high single-photon purity and controlled polarization.
Nevertheless, addressing the issues associated with previously discussed complicated dynamics and emission wobbling of our emitters is crucial for advancing photonic quantum information technology with TMD quantum emitters. These behaviors point to potential sources of additional inhomogeneous broadening, which can become the primary bottleneck for achieving lifetime-limited emission lines, resulting in poor indistinguishability.
Therefore, turning our attention to identifying and dealing with possible inhomogeneous broadening could be the next step toward developing TMD-based quantum emitters with high purity and indistinguishability.
Several strategies could be implemented to deal with such additional broadening.
On the one hand, it is essential to suppress the random charge fluctuations induced by free carriers on the substrate and the direct interaction of the emitters with the environment. One could implement a proper surface stabilization technique, such as encapsulation with AlO via atomic layer deposition Kim et al. (2016) or hBN encapsulation Parto et al. (2021); Kutrowska-Girzycka et al. (2022) which can be directly implemented to the method presented in this work. An alternative approach involves the implementation of electrical contacts Schadler et al. (2019); Zhai et al. (2020). On the other hand, an additional source of emission instability could result from the photo-induced charge carriers under the above-band optical excitation. One way to tackle this limitation lies in the employment of advanced excitation schemes, such as quasi-resonant Bao et al. (2020), resonant Kumar et al. (2016); Somaschi et al. (2016), and SUPER excitation Bracht et al. (2021); Karli et al. (2022); Vannucci et al. (2024). Implementing these schemes Bao et al. (2020); Kumar et al. (2016); Bracht et al. (2021); Karli et al. (2022) in TMD quantum emitters can facilitate the efficient exciton population and the mitigation of additional exciton-phonon interactions.
@@ III Conclusion
In this work, we presented an efficient method for generating highly polarized single photons from orientation-controlled nanowrinkles in monolayer WSe. This method employs novel nanostructures to induce one-directional strain deterministically, enabling efficient control of the polarization of emitted single photons. We have successfully demonstrated single photon emission from WSe nanowrinkle emitters with high purity () and a high degree of polarization (99 4%). These findings provide crucial insights for the future integration of these quantum emitters into a specific mode of desired photonic structures, such as the TE mode of a planar ridge waveguide. Furthermore, understanding the fabrication limitations and critical aspects of deterministic polarization control of quantum emitters are essential steps toward their effective utilization as functional single-photon sources in quantum photonics.
@@ IV.1 Sample preparation
Nanopillar fabrication: The substrate was prepared by thermally oxidizing 110 nm of SiO on a silicon wafer. Alignment marks were patterned by UV lithography (using the AZ 5214E positive resist) and subsequent e-beam evaporation of 5 nm of Ti and 50 nm of Au. After cleaving the wafer into chips, the fabrication of the nanopillars was carried out by spin coating a high-resolution negative e-beam resist (hydrogen silsesquioxane or HSQ, XR-1541-006) at 3000 rpm for 1 min, followed by two soft baking steps at 120C and 220C, both for 2 minutes. The resist was patterned by an e-beam writer (JEOL 9500, 100 kV, 6 nA) with a dose of 11000 µC/cm and developed in a 1:3 solution of AZ 400K : HO to obtain 150 nm-tall glass nanopillars in four different shapes, namely cylindrical pillars, three-pointed stars, five-pointed stars, and bowties. It is worth remarking that a high dose is required to cure the spin-on dielectric, whose composition becomes SiO. Other negative e-beam resists with high sensitivity (activated at doses as low as 30 µC/cm) have also been tested (Supporting Note S5).
@@
Defect fabrication: Defects were introduced on the monolayer lattice by bombarding the strained areas of the material with an e-beam (JEOL 9500 e-beam writer, 100 kV, 1000 µC/cm).
@@ IV.2 Optical characterization
Photoluminescence measurements: The PL measurements were obtained with a custom-built low-temperature micro-photoluminescence (µPL) setup. The sample is mounted on a closed-cycle cryostat (attoDRY800) operating at a base temperature of 4 K. The cryostat is equipped with piezoelectric nanopositioners and a low-temperature microscope objective ( NA = 0.82) located inside the cryostat. The sample was excited with a CW LED at 470 nm for the PL imaging. For the microphotoluminescence spectroscopy, a pulsed (80 MHz) femtosecond laser (Chameleon Ultra II, Coherent) at 532 nm was employed for the optical excitation of the emitters. Further investigation of the effect of the different excitation schemes on the emitters is presented in Supporting Note 3 and Figure S8 of the Supporting Information.
The PL spectra were acquired with a fiber-coupled spectrometer (iHR 550, Horiba). A polarizer mounted on a motorized rotation mount was placed on the collection path for polarization-resolved measurements. PL images were acquired with an integration time of 1 s using a CMOS camera (Tucsen Dhyana, 4 MP).
@@ IV.3 Sample imaging
Figure S1: Schematic representation of the fabrication flow, Figure S2: Comparison of PL spectra before and after e-beam irradiation, Figure S3-5: Investigation on the morphological profile of the nanowrinkles for all different structures, Figure S6: SEM images of TS and FS structures with a statistical analysis on the misalignment of the nanowrinkles, Figure S7: SEM images of BT structures with a histogram of misalignment of the nanowrinkles, Figure S8: Comparison of the PL emission of an exemplary quantum emitter under two excitation schemes (LED, 470 nm and pulsed laser, 532 nm), Figure S9: AFM image of a WSe monolayer flake on a TS structure made from AR-N 7520.11., Figure S10: Comparison between the lifetime of the delocalized exciton and the exemplary SPE studies in Figure 5, Figure S11: Histogram regarding the spectral wobbling of the ZPL studied in Figure 5
@@
The authors acknowledge support from the European Research Council (ERC-StG “TuneTMD”, grant no. 101076437) and the Villum Foundation (grant no. VIL53033). The authors also acknowledge the European Research Council (ERC-CoG “Unity”, grant no. 865230).